OpenStudy (anonymous):

Use integration by parts to solve (x+1)ln(2x)

OpenStudy (anonymous):

\[\int\limits (x+1).\ln (2x)\]

OpenStudy (anonymous):

dv = (x - 1) u = ln(2x) du = 1/x dx v = x^2/2 - x

OpenStudy (anonymous):

So the derivative of ln(2x) is the same as that of ln(x)?

OpenStudy (anonymous):

yea coz its a chain rule and it will be: 1/2x * 2 = 2/2x =1/x

OpenStudy (anonymous):

ahhhh okay. thanks, i just wasn't sure on d/dx of ln(2x)